OpenStudy (anonymous):

Find all the cube roots of 0.000125

OpenStudy (anonymous):

0.000125=125/1000000=(5/100)^3 so what do you think are the answers? @obadkato

OpenStudy (anonymous):

im really not sure -_-

OpenStudy (anonymous):

is this a question of complex number ?@obadkato